OpenStudy (anonymous):

Which expression represents x in the system below? x + 2y = 2 6x – 8y = –3

OpenStudy (anonymous):

do you want to solve for x?

OpenStudy (anonymous):

Both do. Just in different ways.

OpenStudy (anonymous):

Yes I do. I did not mention that sorry

OpenStudy (anonymous):

4(x+2y=2) 6x-8y=-3 10x=5 x=1/2